How long will it take Bonita to save $1,000 for a down payment on a car if she makes
$40 a night babysitting and she typically babysits every Friday and Saturday? Can it be answered ASAP

Answer by math_tutor2020(3816)


Answer: 25 nights
(or 25 days, depending on how you view the problem)

---------------------

Work Shown:
x = number of nights babysitting
1 night = 40*1 = 40 dollars
2 nights = 40*2 = 80 dollars
3 nights = 40*3 = 120 dollars
and so on

In general,
x days = 40x dollars

Meaning we can say
40x = 1000
x = 1000/40
x = 25

Bonita will need to babysit for 25 nights to earn $1000 total.